Circular No. 5904

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET KUSHIDA-MURAMATSU (1993t) S. Nakano, Sumoto, Japan, provides the following preliminary para-

bolic orbital elements from the observations on IAUC 5903 (except Dec. 11). It is possible that this is a short-period comet.

T = 1994 Nov. 26.155 TT  ω = 117.^{o}137
Ω = 86.164 }2000.0
q = 1.13471 AU  i = 10.148
1993/94 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 9 5^{h}35.^{m}54 +22^{o}33.^{′}2 3.639 4.617 172.^{o}6 1.^{o}6 15.9
14 5 31.51 +22 35.2 3.584 4.568 178.5 0.3 15.9
19 5 27.34 +22 36.9 3.537 4.518 175.2 1.0 15.8
24 5 23.12 +22 38.4 3.499 4.469 169.2 2.4 15.7
29 5 18.91 +22 39.6 3.469 4.419 163.2 3.7 15.7
Jan. 3 5 14.80 +22 40.6 3.446 4.369 157.1 5.0 15.6
8 5 10.85 +22 41.6 3.432 4.319 151.1 6.3 15.5
13 5 07.14 +22 42.5 3.424 4.269 145.2 7.6 15.5
18 5 03.72 +22 43.6 3.423 4.218 139.3 8.7 15.4

NOVA CASSIOPEIAE 1993 B. Skiff, Lowell Observatory, reports that a candidate precursor ap-

pears on both prints of the Palomar Sky Survey as a star of B ∼ 18, B–R ∼ +1.0. The candidate is the northern component of a merged pair (separation ∼ 2′′) of roughly equal brightness and elongated north-south. The measured position from the E print, via the Lowell Observatory PDS machine, is α = 23^{h}39^{m}22^{s}.36, δ = +57^{o}14′22′′.9 (equinox 1950.0); a measurement of the O print yields end figures 22^{s}.39, 23′′.2.

Visual magnitude estimates: Dec. 12.50 UT, 6.5 (H. Abe, Yatsuka, Shimane, Japan); 12.77, 6.5 (H. Bengtsson, Hisings Backa, Sweden).

PERIODIC COMET ENCKE Total visual magnitude estimates (cf. IAUC 5896): Dec. 5.14 UT, 12.5

(C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector); 8.02, 13.1: (G. Kronk, Troy, IL, 0.33-m reflector).

1993 December 12 Daniel W. E. Green